OpenStudy (anonymous):

FIRST CORRECT ANSWER GETS A MEDAL!! Which ordered pair maximizes the objective function P = 3x + 8y?

Nnesha (nnesha):

order pair ???

OpenStudy (anonymous):

(0, 0) (2, 7) (5, 6) (8, 1) which one

OpenStudy (stefrheart):

You need to substitute each of those ordered pairs in until you get the largest number so: (0,0)=3(0)+8(0)=0 (2,7)=3(2)+8(7)=62 (5,6)=3(5)+8(6)=63 (8,1)=3(8)+8(1)=42 SO the answer would be (5,6)

OpenStudy (anonymous):

thank you

OpenStudy (stefrheart):

Welcome! :)